In a random sample of 100 units from an assembly line, 15 were defective. Does this constitute sufficient evidence at the 10% significance to conclude that the defective rate among all units exceeds 10%?


1
Expert's answer
2021-03-22T09:55:30-0400

"H_0 : p = 0.1 \\

H_1 : p > 0.1"

If H0 is correct, is correct, "np = 100 \times 0.1=10" and "nq = 100 \times 0.9=90" are both above 5.

The test statistic is The test statistic is Z.

The significance level is 10%, so α = 0.10

This is a right-tail test and the critical value is zα= z0.10 = 1.282

Reject H0 if the observed test statistic is greater than 1.282.

"σ_{\hat{p}}= \sqrt{ \frac{p_0q_0}{n} } = \sqrt{ \frac{0.1 \times 0.9}{100} }= 0.03 \\

z = \frac{\hat{p}-p_0}{σ_{\hat{p}}} = \frac{0.15-0.1}{0.03}=1.667"

z = 1.667 > 1.282 so we can reject H0. At α = 0.10 there is sufficient evidence to conclude that the defective rate is above 10%.